OpenStudy (anonymous):

. A video game store sells new and previously-owned video games. If there are 122 new games and 65 previously-owned games, what is the probability that the next video game purchased is a previously-owned game?

OpenStudy (anonymous):

122+65=x answer is 65/x

OpenStudy (anonymous):

Just like before, the probability of something we care about happening is the number of things we care about divided by the total number of things. In this case, we care about previously owned things. Or, 65/(122+65)=65/187

OpenStudy (anonymous):

65 / 187 right

OpenStudy (anonymous):

ok just checking my work thanks guys

OpenStudy (anonymous):

yup :)